OpenStudy (anonymous):

 sin^2 theta + cos theta =2 (Hint: Use the Pythagorean identity sin^2 theta +cos^2 theta =1 to replace sin^2 theta in the given equation).

OpenStudy (lalaly):

first solve for sin^2theta in the second equation, do you know how?

OpenStudy (anonymous):

0?

OpenStudy (lalaly):

\[\sin^2\theta+\cos^2\theta=1\]therefore\[\sin^2\theta=1-\cos^2\theta\] replacee the sin^2theta in ur first equation by the one i jus found

OpenStudy (anonymous):

so 1-cos^2 theta +cos^2 theta =1

OpenStudy (lalaly):

yes

OpenStudy (anonymous):

so thats my answer?

OpenStudy (lalaly):

but wait you wrote 1-cos^2 theta +cos^2 theta =1 its \[1-\cos^2\theta+\cos \theta =1\]

OpenStudy (lalaly):

if u are asked to find theta, then u just solve for costheta and find the value of theta

OpenStudy (lalaly):

=2*

OpenStudy (lalaly):

\[1-\cos^2\theta+\cos \theta=2\]